OpenStudy (anonymous):

(2^x+1)/(2^x-1)=5 . please solve

Nnesha (nnesha):

\[\huge\rm \frac{2^{x+1}}{2^{x-1}}= 5 \] like this ?

OpenStudy (anonymous):

the 1 is not in the power

OpenStudy (anonymous):

\[(2^x+1)/(2^x-1)=5\]

OpenStudy (dayakar):

\[ 2^{x}=a\] substitute in the problem

OpenStudy (dayakar):

(a+1)/(a-1) = 5

OpenStudy (dayakar):

a+1 =5(a-1)

OpenStudy (dayakar):

a+1 = 5a-5 1+5= 5a-a 6=4a 6/4=a 3/2 =a

OpenStudy (dayakar):

a=3/2 but \[a=2^{x}\]

OpenStudy (dayakar):

\[2^{x}=3/2\]

OpenStudy (dayakar):

is it clear